# Theory of Unsupervised Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Estimating Background Covariances via Super-Resolution

### v4

@@ 6.1 Time evolution of vorticity in the jet system
Figure 9 shows a typical evolution of HR vorticity. Initially ($t=0$), an unstable jet exists around the $y$ center with small perturbations. The jet starts meandering and breaks down into multiple vortices, accompanied by the development of fine filaments ($1\lesssim t\lesssim 4$). Subsequently, vortex merging occurs, intensifying small-scale structures again ($4\lesssim t\lesssim 12$). Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$). This final state is statistically steady and can be understood by Rossby wave dynamics [51].
@@
The maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$. This indicates that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability. Without DA, the deviation from the ground truth continues to increase over time, as shown in Section 6.3.
6.2 Accuracy of SR inference
@@
Flow-field reconstruction has been extensively studied using neural networks [74, 75, 76, 77]. In these studies, inputs are incomplete, such as spatially sparse sensor measurements. Neural networks are then trained by comparing their outputs to complete target data. Our problem setup is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
The success of SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution. The loss function $l_{\text{SR}}$ in Eq. (18) operates at individual grid points, and only the non-missing observed points contribute to $l_{\text{SR}}$. This point-wise nature enables backpropagation even when the target data are missing. Furthermore, since convolution shares its kernel across space (known as weight sharing) [78], the efficient learning was realized using the small amount of non-missing observational data.
@@
The trained SR model makes non-local inference. This non-locality is essential for assimilating observations, as discussed in Section 4.3. We use two indicators for interpreting the SR model: gradients and local attribution maps (LAMs) [31]. Gradients are expressed as $\partial\mathsf{F}(\bm{x})/\partial\bm{x}$, where $\mathsf{F}$ is the trained SR model. Such a simple gradient is sometimes not a good indicator for interpreting how a model output depends on its input due to nonlinearity [31, 79]. LAM, proposed by Gu and Dong [31], combines gradient methods with path integral techniques for interpreting behaviors of nonlinear SR models.
Figure 11 shows gradient and LAM fields for the three SR samples previously discussed in Fig. 10. We focus on the square regions (black boxes) in the figures. The gradients and LAMs indicate that colored points in the LR inputs contribute to the HR outputs in these boxes. If there are observed values in these boxes, the colored points in the inputs are updated by minimizing the loss function (Section 4.3). Thus, observations have a non-local influence through the trained SR model. This is an essential reason for the success of the proposed SRDA. Interestingly, the LAMs tend to take larger values over wider regions, compared to the gradients (Fig. 11), suggesting the nonlinearity of the SR model. For instance, in the first row of the figure, the LAM has a larger value at one grid point below the black box, matching the position of the vortex filament.
Figure 11: Attribution maps for SR. The left two columns, SR output and LR input, are the same as in Fig. 10, but shown here for comparison. Local attribution maps (LAMs) [31] were calculated for the small square regions of $11\times 1$ 11 (i.e., black boxes in the figure). These boxes were specified on the HR grids, so they are not necessarily aligned with the LR grids. The gradients and LAMs were normalized by their maximum values, with ranges between 0 and 1, following Gu and Dong [31].
6.3 Accuracy of LR forecast
@@
Figure 13 shows the time series of the mean values for the MAE ratio and MSSIM loss. Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (Section 6.1). In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines). The error reduction occurs every time unit since the assimilation interval was set to 1. When smoothed over time, the time series of the EnKF and SRDA have two peaks, corresponding to the two developments of small-scale patterns during the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). The bottom panels in Fig. 13 show the differences in the test errors between the EnKF and SRDA, where the gray areas represent the 10th to 90th percentile values. The test errors of the SRDA are significantly smaller than those of the EnKF. Thus, we conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
@@
In the proposed SRDA, the LR forecast can be super-resolved to the HR at an arbitrary time using the SR model (Section 4.6). Figure 14 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast. The merging of two vortices is not obvious in the LR, but is well captured in the super-resolved forecast.
@@ 6.4 Accuracy of HR analysis
Figure 17 compares the analyses and their SDs for the EnKF and SRDA. The SRDA SD tends to be large around vortex filaments or vortex edges. This tendency appears to be most evident during the jet collapse ($t=2$). This result is reasonable because vortex locations are generally uncertain, and small changes in their locations lead to large differences in vorticity. The EnKF SD also exhibits a similar tendency; however, the SD distribution is blurred, likely reflecting the blurred vortex patterns. The SD value ranges also differ: the EnKF SD is between 0.3 and 0.9, while the SRDA SD is between 0.8 and 0.9. Compared to the EnKF, the SRDA tends to estimate larger SD in regions where vorticity is close to zero, especially at $t=9$ and 11 in Fig. 17. Although the true value of SD is unknown, our results suggest that the SRDA tends to estimate larger SD in low-vorticity regions. This tendency may be due to a simplified variance estimation in the present CVAE (Section 5.3), where the variances are computed using the SR model from the LR analysis. The SRDA demonstrated a potential for uncertainty quantification without relying on ensemble calculation; however, further model development is necessary. Recent studies demonstrate that VAEs can learn error distributions of model forecasts [22, 23, 24, 25]. Their model architectures may be effective for the SRDA.
@@
6.5 Computation time of SRDA
@@
7 Conclusions
This study has proposed a theory of unsupervised SRDA with the CVAE. Using the ELBO, we have developed an unsupervised theory of DA, which is an extension of the 3D-Var (Section 3). Since ELBOs generally serve as the loss function for CVAEs, the assimilation is feasible using the CVAE. In this theory, it is essential to use a non-diagonal background error covariance matrix, as in the 3D-Var. Instead of using such non-diagonal matrices, we have extended the theory by using the non-locality of SR to perform DA (Section 4). This extension provides a theoretical foundation for combining SR and DA. The combination of SR models and CVAEs enables us to assimilate HR observations into LR forecasts computed with LR fluid models. This approach eliminates the need for time integration of HR fluid models, thereby reducing computational burden.
The effectiveness of the proposed method was evaluated via numerical experiments using an idealized barotropic ocean jet system (Sections 5 and 6). Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times. Our results suggest the potential of SRDA as a promising approach for integrating observations into CFD simulations while obtaining HR inferences with a low computational cost.
There are several future research topics for SRDA. The proof of concept here has only examined the linear observation operator, although the proposed SRDA theoretically allows learning of nonlinear observation operators. Furthermore, advanced training methods, such as online training [63, 64] and loss balancing for VAEs [68], will be important to develop SRDA. As for generalizability, we have not evaluated to what extent the CVAE remains effective when the statistical properties of the input data differ. These aspects should be addressed in a more realistic experimental setup. Indeed, the fluid system used here is simple compared to actual atmospheric and oceanic systems. It is necessary to evaluate the proposed SRDA for more complex systems.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
@@
Appendix A Hyperparameters of the SR model and the CVAE
We list the hyperparameters of the SR model and the CVAE shown in Fig. 7. Table 2 shows the numbers of channels and strides in all convolution layers. The kernel size of all convolutions is 3. The slope of all leaky ReLU [59] is $-0.01$. The upscale factor of all pixel shuffle operations [57] is 2. The resizing in the SR model is performed by bicubic interpolation. The weight parameters of all convolution layers are initialized randomly using uniform distributions [80].
@@ Appendix A Hyperparameters of the SR model and the CVAE
Data availability
The data and source code that support the findings of this study are preserved and openly available at our Zenodo repository (https://doi.org/10.5281/zenodo.13958042). These files are also available at our GitHub repository (https://github.com/YukiYasuda2718/srda-cvae/releases/tag/v0.2.1).
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used Claude 3.5 only for English editing. After using this service, the authors reviewed and edited the content as needed. The authors take full responsibility for the content of the publication.

### v5

@@ VI.1 Time evolution of vorticity in the jet system
Figure 9 shows a typical evolution of HR vorticity. Initially ($t=0$), an unstable jet exists around the $y$ center with small perturbations. The jet starts meandering and breaks down into multiple vortices, accompanied by the development of fine filaments ($1\lesssim t\lesssim 4$). Subsequently, vortex merging occurs, intensifying small-scale structures again ($4\lesssim t\lesssim 12$). Finally, a wavenumber-1 structure continues to propagate westward ($12\lesssim t$). This final state is statistically steady and can be understood by Rossby wave dynamics.David, Marshall, and Zanna (2017)
@@
The maximum Lyapunov exponent $\lambda$ is positive: $\lambda=3.11$. This indicates that the temporal evolution of vorticity exhibits initial-value sensitivity, likely due to the barotropic instability. Without DA, the deviation from the ground truth continues to increase over time, as shown in Section VI.3.
VI.2 Accuracy of SR inference
@@
Flow-field reconstruction has been extensively studied using neural networks.Lluís et al. (2020); Maulik et al. (2020); Gundersen et al. (2021); Wang et al. (2022) In these studies, inputs are incomplete, such as spatially sparse sensor measurements. Neural networks are then trained by comparing their outputs to complete target data. Our problem setup is opposite to that of these studies: in our case, the input is complete, while the target (i.e., the observation) is incomplete.
The success of SR in our study is attributable to two factors: (i) the point-wise nature of the loss function and (ii) the weight sharing of convolution. The loss function $l_{\text{SR}}$ in Eq. (18) operates at individual grid points, and only the non-missing observed points contribute to $l_{\text{SR}}$. This point-wise nature enables backpropagation even when the target data are missing. Furthermore, since convolution shares its kernel across space (known as weight sharing),Aloysius and Geetha (2017) the efficient learning was realized using the small amount of non-missing observational data.
@@
The trained SR model makes non-local inference. This non-locality is essential for assimilating observations, as discussed in Section IV.3. We use two indicators for interpreting the SR model: gradients and local attribution maps (LAMs).Gu and Dong (2021) Gradients are expressed as $\partial\mathsf{F}(\bm{x})/\partial\bm{x}$, where $\mathsf{F}$ is the trained SR model. Such a simple gradient is sometimes not a good indicator for interpreting how a model output depends on its input due to nonlinearity.Gu and Dong (2021); Sundararajan, Taly, and Yan (2017) LAM, proposed by Gu and Dong,Gu and Dong (2021) combines gradient methods with path integral techniques for interpreting behaviors of nonlinear SR models.
Figure 11 shows gradient and LAM fields for the three SR samples previously discussed in Fig. 10. We focus on the square regions (black boxes) in the figures. The gradients and LAMs indicate that colored points in the LR inputs contribute to the HR outputs in these boxes. If there are observed values in these boxes, the colored points in the inputs are updated by minimizing the loss function (Section IV.3). Thus, observations have a non-local influence through the trained SR model. This is an essential reason for the success of the proposed SRDA. Interestingly, the LAMs tend to take larger values over wider regions, compared to the gradients (Fig. 11), suggesting the nonlinearity of the SR model. For instance, in the first row of the figure, the LAM has a larger value at one grid point below the black box, matching the position of the vortex filament.
Figure 11: Attribution maps for SR. The left two columns, SR output and LR input, are the same as in Fig. 10, but shown here for comparison. Local attribution maps (LAMs)Gu and Dong (2021) were calculated for the small square regions of $11\times 1$ 11 (i.e., black boxes in the figure). These boxes were specified on the HR grids, so they are not necessarily aligned with the LR grids. The gradients and LAMs were normalized by their maximum values, with ranges between 0 and 1, following Gu and Dong.Gu and Dong (2021)
VI.3 Accuracy of LR forecast
@@
Figure 13 shows the time series of the mean values for the MAE ratio and MSSIM loss. Without SR or DA (green dotted lines), the test errors increase over time due to the sensitivity to initial conditions (Section VI.1). In contrast, this error growth is suppressed by the EnKF (orange dashed lines) and the SRDA (blue solid lines). The error reduction occurs every time unit since the assimilation interval was set to 1. When smoothed over time, the time series of the EnKF and SRDA have two peaks, corresponding to the two developments of small-scale patterns during the jet collapse ($t\sim 2$) and the vortex merging ($t\sim 9$). The bottom panels in Fig. 13 show the differences in the test errors between the EnKF and SRDA, where the gray areas represent the 10th to 90th percentile values. The test errors of the SRDA are significantly smaller than those of the EnKF. Thus, we conclude that the SRDA forecast outperforms the EnKF in terms of pixel-wise accuracy and pattern consistency.
@@
In the proposed SRDA, the LR forecast can be super-resolved to the HR at an arbitrary time using the SR model (Section IV.6). Figure 14 shows a typical vortex merging between $t=8.5$ and 11.0 from the ground truth, the LR forecast, and the super-resolved forecast. The merging of two vortices is not obvious in the LR, but is well captured in the super-resolved forecast.
@@ VI.4 Accuracy of HR analysis
Figure 17 compares the analyses and their SDs for the EnKF and SRDA. The SRDA SD tends to be large around vortex filaments or vortex edges. This tendency appears to be most evident during the jet collapse ($t=2$). This result is reasonable because vortex locations are generally uncertain, and small changes in their locations lead to large differences in vorticity. The EnKF SD also exhibits a similar tendency; however, the SD distribution is blurred, likely reflecting the blurred vortex patterns. The SD value ranges also differ: the EnKF SD is between 0.3 and 0.9, while the SRDA SD is between 0.8 and 0.9. Compared to the EnKF, the SRDA tends to estimate larger SD in regions where vorticity is close to zero, especially at $t=9$ and 11 in Fig. 17. Although the true value of SD is unknown, our results suggest that the SRDA tends to estimate larger SD in low-vorticity regions. This tendency may be due to a simplified variance estimation in the present CVAE (Section V.3), where the variances are computed using the SR model from the LR analysis. The SRDA demonstrated a potential for uncertainty quantification without relying on ensemble calculation; however, further model development is necessary. Recent studies demonstrate that VAEs can learn error distributions of model forecasts.Grooms (2021); Yang and Grooms (2021); Grooms et al. (2023); Xiao et al. (2025) Their model architectures may be effective for the SRDA.
@@
VI.5 Computation time of SRDA
@@
VII Conclusions
This study proposed a theory of unsupervised SRDA using the CVAE. By deriving an ELBO for DA, we showed that the theory is an extension of the 3D-Var (Section III). Since ELBOs generally serve as the loss function for CVAEs, the assimilation is feasible using the CVAE. In this theory, it is essential to use the non-local SR operation, which makes background covariances learnable from data (Section IV). The proposed SRDA assimilates HR observations into LR forecasts without running HR fluid models. The effectiveness of the proposed method was evaluated via numerical experiments using an idealized barotropic ocean jet system (Sections V and VI). Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times.
There are several future research directions for unsupervised SRDA. The major limitation of the proposed theory is the Gaussian assumption Eq. (10), which suggests that the accuracy in inference would be reduced if a system exhibited strong non-Gaussianity. One method to extend the theory to non-Gaussian systems is combining CVAEs with normalizing flowsRezende and Mohamed (2015); Kobyzev, Prince, and Brubaker (2021) to learn non-linear transformations that lead to non-Gaussian distributions. The proof of concept here has only examined the linear observation operator, although the proposed SRDA theoretically allows learning of nonlinear observation operators. Furthermore, advanced training methods, such as online trainingSahoo et al. (2018); Rasp (2020) and loss balancing for VAEs,Asperti and Trentin (2020) will be important to develop SRDA. As for generalizability, we have not evaluated to what extent the CVAE remains effective when the statistical properties of the input data differ. These aspects should be addressed in a more realistic experimental setup. Indeed, the fluid system used here is simple compared to actual atmospheric and oceanic systems. It is necessary to evaluate the proposed SRDA for more complex systems.
Acknowledgements.
@@
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose
Data Availability Statement
The data and source code that support the findings of this study are openly available at our Zenodo repository (https://doi.org/10.5281/zenodo.13958042), reference number Yasuda and Onishi, 2024. These files are also available at our GitHub repository (https://github.com/YukiYasuda2718/srda-cvae/releases/tag/v0.2.1).
Appendix A Hyperparameters for the SR model and the CVAE
We list the hyperparameters of the SR model and the CVAE shown in Fig. 7. Table 2 shows the numbers of channels and strides in all convolution layers. The kernel size of all convolutions is 3. The slope of all leaky ReLUMaas et al. (2013) is $-0.01$. The upscale factor of all pixel shuffle operationsShi et al. (2016) is 2. The resizing in the SR model is performed by bicubic interpolation. The weight parameters of all convolution layers are initialized randomly using uniform distributions.He et al. (2015) For the Adam optimizer, the default parametersKingma and Ba (2015) are used except for the learning rate of $1\times 10^{-4}$. The mini-batch size is set to 128 for all experiments.